# Light-induced ultrafast magnetization dynamics in van der Waals antiferromagnetic CrSBr

### v2

@@ Abstract
The magnetization dynamics driven by the femtosecond laser pulse of antiferromagnet van der Waals semiconductor CrSBr is studied within time-dependent density functional theory. We investigate the effect of laser fluence as well as the excitation frequency on the ultrafast dynamics of spins. In low fluence, the local magnetic moment of Cr increases when the laser frequency is below the band gap, whereas it decreases when the laser frequency is above the band gap. In high fluence, we find strong demagnetization independent of excitation frequency. We find that the ultrafast demagnetization in CrSBr is dominated by intralayer and interlayer optical intersite spin transfer, and spin flip via the spin-orbit coupling plays a minor role. Our results reveal hole excitation in the low-fluence regime and electron excitation in the high-fluence regime. Further, we investigate the effect of the external static magnetic field on the dynamics. We demonstrate even and odd high harmonic generation in the local magnetic moments and electric current, respectively. The external magnetic field results in an out-of-plane charge current via the inverse spin Hall effect as well as odd harmonics in the local magnetic moment dynamics as a result of breaking the time-reversal symmetry.
Introduction.— Van der Waals (vdW) magnets provide a promising platform for applications in both quantum and classic information processing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. They exhibit atomically sharp interfaces enabling heterostructures engineering [11] as well as exceptional tunability through electric [12, 13, 14], magnetic [15, 16], mechanical [17, 18, 19, 18], and optical means [17, 9]. Particularly, vdW magnets can exhibit semiconducting electronic structure, facilitating strong couplings between magnetic structures and optical excitations [20, 15, 16, 21]. CrSBr is a prototype antiferromagnetic semiconductor showing strong coupling between microwave-range magnon and optical-range excitons [15, 16, 22], which provides an ideal platform for bridging microwave and optical regimes [23, 24, 25], positioning as a new frontier for spintronics, magnonics, and quantum information technologies.
Understanding and controlling vdW magnetic materials are crucial [26, 27, 28]. Ultrafast laser pulses have become a crucial and powerful means since the observation of the pulse-induced demagnetization in nickel within a sub-picosecond time scale [29], providing an ultrafast and non-thermal way to probe and control magnetic materials [28]. It has been shown that, when CrSBr is excited by ultrafast laser pulse, intriguing phenomena like periodically modulated exciton energy, nonlinear magnon frequency mixing, fast spin wavepacket transport, and even extreme high harmonic generations [15, 16, 30, 31, 32, 33], have been observed.
However, a microscopic understanding of how coherent spin dynamics is excited is still missing, especially when the nonequilibrium optical excitations are concerned, where magnetization from the excitations and local -orbital electrons are entangled. In this work, we present a full ab initio description of the ultrafast magnetization dynamics in CrSBr using real-time TDDFT (rt-TDDFT). Indeed, TDDFT gives a microscopic understanding of the ultrafast demagnetization, when the magnetization amplitude changes and the phenomenological Landau-Lifshitz-Gilbert formalism of magnetization with fixed amplitude, is not valid anymore. Here, we use TDDFT to systematically study the effect of the laser fluence, frequency, and magnetic fields on the magnetization dynamics of the local atomic moments in CrSBr. While the previous experimental and theoretical studies always showed demagnetization of the local moments in magnetic metals, here we observe an increase in these moments when the system is excited with a low fluence laser pulse with below gap central frequency. With low fluence laser, we find that the local moments show strong dependence on the laser central frequency. At high fluence, we find demagnetization of the local magnetic moments that is independent of the laser frequency.
Formalism.— CrSBr is an A-type vdW antiferromagnetic semiconductor where Cr atoms have nonzero local magnetic moment. The intralayer and interlayer magnetic orders are ferromagnetic and antiferromagnetic, respectively, with triaxial magnetic anisotropy and the easy axis being in the plane, Fig. 1 (a) [34]. It is air stable material with bulk Néel temperature K [22].
@@
We employ the start-of-the-art rt-TDDFT method to study the ultrafast magnetization dynamics in bulk CrSBr excited by a laser pulse. The effect of laser pulse is captured by the time-dependent Kohn-Sham equation for spinor Kohn-Sham orbitals (using )
@@
where is the vector potential of the laser pulse; is the speed of light; is the effective time-dependent Kohn-Sham potential, contributed by external potential from nuclei , Hartree potential , and exchange potential ; is the time-dependent Kohn-Sham magnetic field with being the external magnetic field and the exchange magnetic field; are Pauli matrices; and the last term describes spin-orbit coupling.
We employ adiabatic local spin density approximation (ALSDA) for exchange functional within full-potential augmented plane-wave method as implemented in the ELK code [35, 36]. The calculation is done in two stages; first, the ground state is obtained from ELK using static DFT calculations with local density approximation (LDA) exchange functional. We use DFT+U with eV, eV, and fully localized limit double counting scheme is applied. The ground-state band gap is eV and the magnetic moment for Cr atoms in agreement with literature [37, 38]. Next, the time-dependent calculation is performed using the ground-state Hamiltonian of the first step with a Gaussian femto-second laser pulse. In this study, the laser pulse is polarized along -direction traveling along -direction with wavelength of 1240 nm (below the band gap (), ) and 620 nm (above the band gap, ), full width at half maximum of 12 fs, and different laser pulse energies. The simulation time is 150 fs and a time step of 0.6 attosecond is chosen. The grid of vectors is chosen as . Since the wavelength of applied laser light is much larger than the supercell, we assume dipole approximation and disregard spatial dependence of the vector potential . In calculations with the external magnetic field, a static external magnetic field () is added to the Hamiltonian in both ground-state and time-dependent simulations.
@@
where is the total potential within the crystal, is the charge current along the -direction, and is the spin current along the -direction and -spin polarization, . Equation (2) demonstrates that magnetization dynamics is driven not only by external fields, but also by spin currents from optically excited carriers as well as spin-orbit coupling that is manifested as spatial derivatives of the potential. In this study, the ions are assumed fixed in the time-dependent calculations and we only focus on about 100 fs after the optical excitation where the most relevant dynamics is driven by the electronic system and this approximation is to a good extend valid.
@@
Low fluence excitation.— Here we study the effect of the low-fluence () light on the magnetization dynamics. We find that the magnetization change shows a strong dependence on the light frequency. When the light frequency is below the band gap, the local magnetic moments of the Cr atoms are enhanced as shown in Fig. 1(a). This figure displays the time evolution of the magnetic moment of the Cr atom () in the layer that the magnetization points along the direction (which lies along the lattice vector ). The magnetic moments of Cr atoms in the other layer are opposite leading to zero net magnetization, guaranteed by the two-fold rotation symmetry about -axis. The laser pulse is illustrated by the black curve that lasts for about 12 fs. After the pulse, increases by about 2%. The blue solid line and red dotted line show the calculations with and without SOC, respectively. One can find that the effect of SOC is weak and does not change the qualitative behavior of the evolution of .
@@
The increase of the magnetic moment arises from spin transfer between magnetic and nonmagnetic atoms. As shown in Fig. 1(b) calculated with SOC, the charge density of majority spin in Cr atom increases while that in S and Br atoms in the same layer decreases. For spin minority in Fig. 1(c), the trend is opposite. This indicates net spin transfer from S and Br to Cr, increasing the magnetic moment of the Cr atoms. In addition, we note that the reduction of spin in intralayer S and Br atoms is smaller than the increase of spin in Cr atom, suggesting that there is interlayer spin transfer. The spin transfer process is analogous to that in metallic materials where both magnetic and nonmagnetic atoms are present. However, in those systems, the spin transfer, termed optical inter-site spin transfer (OISTR) [40], typically induces demagnetization. Here we show that in the antiferromagnetic semiconductor, the OISTR can also increase the magnetic moment of magnetic atoms.
@@
The spin transfer shows strong dependence on the light frequency. When the light frequency is above the band gap, the local magnetic moments of the Cr atoms are reduced or so-called demagnetized as shown in Fig. 1(e). Similar to the excitation with the below-gap frequency, the effect of SOC is weak, which slightly enhances demagnetization. The mechanism of the change of magnetic moment in Cr atom is similar. Figures 1(f) and 1(g) show that the averaged atomic spin density of S and Br decreases, whereas the spin density of Cr first increases and then reduces to amounts less than the equilibrium in the majority spin channel. In addition, the intralayer spin increase of Br and S is much weaker than the spin decrease in Cr atom, indicating that the interlayer spin transfer dominants the demagnetization.
High fluence excitation.— In the high fluence case (), we find that the qualitative behavior of spin dynamics shows weak dependence on the laser frequency. We focus on the case with below-gap frequency. With collinear antiferromagnetic order, the laser pulse does not induce a net magnetization. The magnetizations in neighboring layers are still opposite. Figure 2(a) plots the magnetic moment of the Cr atoms in one sublattice, where we find strong demagnetization on this sublattice in the presence of a laser pulse. One can see that the dynamics of starts with a slight enhancement, and then after a few fs, the demagnetization initiates. We observe about 50% demagnetization, in contrast to the low fluence cases, which shows weak dependence on the SOC. The effects of SOC start to manifest themselves near the end of the laser pulse, which introduces a weak correction that does not change the qualitative behavior.
The demagnetization in Cr atom leads to a reduction in the Néel vector despite a slight increase in the total magnetic moment of S and Br as shown in Figs. 2(b) and 2(c). In the absence of SOC, the spin flip scattering is forbidden. Therefore, the presence of demagnetization in one layer indicates the presence of interlayer spin transfer. In the presence of SOC, the demagnetization is slightly enhanced, indicating that the SOC-induced spin flip scattering can enhance the demagnetization.
Our further analysis of the time-dependent atomic spin densities are shown in Fig. 2 (b), (c), (f), and (g). Form Fig. 2 (b) and (f) one can see that the symmetric phase of the dynamics actually enhances the magnetic moments of the Cr atoms and this is the reason of the slight enhancement of the that was already discussed. However, the interlayer OISTR very soon dominates the dynamics resulting in fast demagnetization of . On the other hand, Fig. 2 (c) and (g) show a more lasting symmetric phase and therefore a stronger intralayer OISTR which results in demagnetization. Later on, the interlayer OISTR participates which is evident from the the nonsymmetrical phase of the dynamics. Here again, one may conclude that the interlayer and intralayer OISTR mechanisms are highly dependent on the spin channel and the excitation frequency. Spin-flip scatterings via SOC plays a small portion of the demagnetization process and comes into play with a delay.
Free carrier density.—Fig. 3 shows the densities of the excited free carriers i.e. the charge density in the interstitial region for both high and low fluence excitations with AG and BG frequencies. As can be seen in Fig. 3(a) in the low fluence excitation, the excitation results in hole generation while in the high fluence excitation, the excitation results in electron generation after the laser pulse is gone as can be seen in Fig. 3(b). This does not depend on the excitation frequency which is rather surprising in the low fluence case regarding the strong frequency dependence of the atomic spin densities and local magnetic moment of Cr atoms. In both figures, one can see a much larger amplitude of the variation in the case of BG excitation. This is because the AG and BG excitations are dominated by resonant and non-resonant transitions, respectively [41].
Effect of external magnetic field.— In the presence of the external magnetic field, the dynamics imply the generation of transverse magnons. The dynamics of the magnetic moment of Cr atoms along the x and z directions and in the presence of an external magnetic field are shown in Fig. 1 (d) and (h) for the AG and BG excitations, respectively. We note that the dynamics along the y direction in the presence of the external magnetic field is very similar to the dynamics in the absence of the external magnetic field shown in Fig. 1 (a) and therefore are not shown for brevity. We observe the initiation of low frequency oscillations of the magnetic moments along x and z direction in the presence of the external magnetic field indicating magnon generation. In addition, while the dynamics are similar for AG and BG excitations in the absence of the external magnetic field, they are quite different in the presence of the external magnetic field. The same conclusion can be obtained for the high fluence excitation, i.e. external magnetic field causes transverse magnon generation when the system is optically excited.
Electron dynamics and HHG.— In Fig. 4, we plotted the Fourier transformation of the charge current along the y-direction (the other two components are zero) as well as the magnetization of a typical Cr atom along the y-direction. As can be seen from Fig. 4(a), (b), (d), and (e), the response of CrSBr to the external laser pulse excitation contains odd harmonics in the electric current and as a consequence of inversion symmetry the even harmonics are canceled out. In all cases, high harmonics up to 7th harmonics are pronounced. While the higher harmonics are reduced in high fluence excitation, one can observe that the third harmonics survives when the excitation frequency is BG. In Fig. 4(c), the magnetization only contains the even harmonics. This is a consequence of SU(2) symmetry of the spin configuration along this direction [42] and can be seen from the magnetization dynamics in real time in Fig. 1(a) and (e) as well as Fig. 2(a) and (e). Nonetheless, the magnetization dynamics along the other directions (x and z) contain all even and odd harmonics (not shown here) which is driven by SOC interaction. However, the external magnetic field results in the breaking of the time-reversal symmetry and therefore the magnetization contains odd harmonics as well in this case (Fig. 4(f)). It should be noted that in this case, the z components of the electric current becomes nonzero due to the inverse spin Hall effect. In other word, the spin currents driven by the laser pulse, in the presence of the static external magnetic field, generate a perpendicular charge current along the z-direction.
Conclusion.—In this study, the ultrafast dynamics of semiconducting antiferromagnetic CrSBr is investigated using rt-TDDFT formalism. The ground state of the system was excited using a fs laser pulse with low and high fluence and below and above-gap central laser frequencies. We found high fluence excitation results in considerable demagnetization of the local magnetic moment of Cr atoms and it is almost independent of frequency. On the other hand, the dynamics of the magnetic moments of the Cr atoms strongly depends on the frequency in the low fluence excitation resulting in a slight demagnetization of Cr atoms with AG excitation. On the other hand, BG excitation enhances the magnetic moment of the Cr atoms; a peculiar observation which is just recently reported and is not yet carefully studied. In addition, an external applied magnetic field during the optical excitation triggers magnon generation in the nonequilibrium system. The microscopic mechanism of the demagnetization are found to be mostly intralayer and interlayer OISTR and SOC plays minor role. We observe hole generation in low fluence and electron generation in the high fluence excitation with both AG and BG frequencies. In addition, we observe odd and even harmonics generations in the charge current and magnetization, respectively.
Note added— Upon completing this work, we noticed a similar study the monolayer CrSBr [43].

### v3

@@ Abstract
We investigate the ultrafast magnetization dynamics of semiconducting antiferromagnetic CrSBr using real-time time-dependent density functional theory. In zero magnetic field, laser excitation modifies only the magnetization along the easy axis, leaving transverse components unaffected. We find that below-gap, low-fluence pulses enhance the local magnetic moments via spin transfer from nonmagnetic to magnetic atoms, increasing the Néel vector. In contrast, high-fluence pulses drive interlayer spin transfer between magnetic atoms, producing strong demagnetization and reducing the Néel vector, while S and Br atoms exhibit primarily charge transfer with weak opposite contribution to the demagnetization. An applied magnetic field qualitatively alters the response, enabling both magnitude changes and ultrafast reorientation of the magnetization. We show that the resulting layer-resolved reorientation respects a twofold rotation about the -axis, exciting coherent optical magnons even under this symmetry, which modulate the relative angle between neighboring layers and periodically tune electronic properties. These results reveal a microscopic pathway for coherent magnon excitation in van der Waals magnets and establish a framework for controlling their coupled spin–charge dynamics on femtosecond timescales.
Controlling magnetization and its dynamics is central to both the fundamental understanding of magnetic materials and their applications in spintronics and magnonics [1, 2, 3, 4]. The intrinsic timescale of magnetization dynamics is set by the exchange energy, typically lying in the gigahertz to terahertz range [1, 2]. Exciting magnetic materials with ultrafast laser pulses provides a means to bypass this limitation, enabling magnetization control on femtosecond to picosecond timescales [5, 6, 7, 8, 9]. In the past two decades, significant progress has been made in ultrafast control of the magnetic order parameter [6, 7, 8, 9], focusing on the demagnetization processes in transition metals, such as the ultrafast demagnetization in ferromagnets like Ni [6], switching from uncompensated antiferromagnets to transient ferromagnets in superstructures like Co/Mn multilayers[10, 11], and strong terahertz emission associated with the ultrafast demagnetization [12, 13].
The discovery and study of van der Waals magnetic materials have expanded rapidly in recent years showing diverse magnetic order and rich electronic structures, such as metals and semiconductors [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Ultrafast lasers provide powerful means to characterize and manipulate magnetic order and dynamics [24, 25, 26, 27] surpassing other techniques such as electric [28, 29, 30], magnetic [31, 32], mechanical [27, 33, 34, 33] in speed. In particular, experiments have shown that ultrafast laser pulses can serve as an efficient, nonthermal method for generating coherent magnetization dynamics, which in turn dynamically and periodically modulate the electronic and optical properties of these quantum materials [35, 36, 31, 32, 37]. While phenomenological models provide intuitive descriptions of long-lived coherent dynamics [38], the microscopic origin of ultrafast magnetization dynamics—its initiation and its interplay with demagnetization—remains poorly understood [39, 40, 41, 42, 43, 44, 45, 46].
In this work, we present a fully ab initio description of ultrafast magnetization dynamics in a van der Waals magnet using real-time time-dependent density functional theory (rt-TDDFT). Our study focuses on CrSBr, a prototype semiconducting antiferromagnet showing strong coupling between magnetization and electronic excitations, as well as its large-amplitude, laser-pulse-induced magnetization dynamics [47, 31, 32]. We systematically study the effect of the laser fluence, frequency, and magnetic fields. We show that, while the previous experimental and theoretical studies always showed demagnetization of the local moments in magnetic metals [48, 49], our results suggest an increase in the magnetic moments when the system is excited with a low fluence laser pulse with below gap central frequency. Above-gap or high-fluence laser induces demagnetization. In the presence of magnetic fields, we identify ultrafast reorientation of magnetization vectors that excite coherent magnon dynamics in sub-picosecond, despite that the magnon frequencies are in the order of gigahertz. These results deepen our understanding of how the coherent magnons are excited in van der Waals magnets, offering a solid starting point for future study of the subsequent magnetization dynamics.
@@
Formalism.— We employ the start-of-the-art rt-TDDFT method to study the ultrafast magnetization dynamics excited by a laser pulse. The effect of laser pulse is captured by the time-dependent Kohn-Sham equation for Kohn-Sham orbitals as Pauli spinors (using )
@@
where is the momentum operator, is the vector potential of the laser pulse, and is the speed of light. is the effective time-dependent Kohn-Sham potential contributed by external potential from nuclei , Hartree potential , and exchange potential . is the time-dependent Kohn-Sham magnetic field with the external static magnetic field and the exchange magnetic field. are Pauli matrices. The last term describes spin-orbit coupling.
@@
where is the spin current with spin polarized along the -direction with the local current density operator. Equation (2) demonstrates that magnetization dynamics is driven not only by external fields, but also by spin currents from optically excited carriers as well as spin-orbit coupling that is manifested as spatial derivatives of the potential. In this study, the ions are assumed fixed in the time-dependent calculations and we only focus on about 100 fs after the optical excitation where the most relevant dynamics is driven by the electronic system and this approximation is to a good extend valid.
We employ adiabatic local spin density approximation (ALSDA) for exchange functional within full-potential augmented plane-wave method as implemented in the ELK code [51, 52]. The calculation is done in two stages. First, the ground state is obtained using static noncollinear DFT calculations. CrSBr is an A-type vdW antiferromagnetic semiconductor. The intralayer magnetic order is ferromagnetic with two Cr atoms, e.g., Cr1 and Cr2 in Fig. 3(a), in one unit cell. The interlayer magnetic order is antiferromagnetic where two magnetic layers with opposite spin orientations are labeled as layer 1 (L1) containing Cr1 and Cr2 and layer 2 (L2) containing Cr3 and Cr4. CrSBr has a triaxial magnetic anisotropy with easy axis along -direction [53, 47]. We use DFT+U with eV, eV, and fully localized limit double counting scheme is applied. The ground-state band gap is eV and the magnetic moment for Cr atoms in agreement with experimental studies [54, 55]. Throughout the manuscript, , , and directions are along , , and -axis, respectively [53, 47]. Next, the time-dependent calculation is performed using the ground-state Hamiltonian of the first step with a Gaussian femto-second laser pulse. In this study, the laser pulse is polarized along -direction traveling along -direction with wavelength of 1240 nm (below the band gap , ) and 620 nm (above the band gap, ), full width at half maximum of 12 fs, and different laser pulse energies. The simulation time is 150 fs and a time step of 0.6 attosecond is chosen. The grid of vectors is chosen as . Since the wavelength of applied laser light is much larger than the supercell, we assume dipole approximation and disregard spatial dependence of the vector potential . In calculations with the external magnetic field, a static external magnetic field () is added to the Hamiltonian in both ground-state and time-dependent simulations.
@@
Low fluence excitation.— At low fluence (mJ/cm2), we find that the magnetization change shows a strong dependence on the laser frequency. When the laser frequency is below the band gap, we find that the local magnetic moments of the Cr atoms are enhanced. Fig. 1(a) displays the time evolution of the magnetic moment of the Cr atom () in one of the magnetic sublattices along the direction (the easy axis). The magnetic moments of Cr atoms in the other layer are opposite leading to zero net magnetization, guaranteed by the two-fold rotation symmetry about -axis. The laser pulse is illustrated by the black curve that lasts for about 12 fs. After the pulse, increases by about 2%. The blue solid line and red dotted line show the calculations with and without SOC, respectively. One can find that the effect of SOC is weak and does not change the qualitative behavior of the evolution of .
@@
The increase of the magnetic moment arises from spin transfer between magnetic and nonmagnetic atoms. As shown in Figs. 1(b) and 1(c), the charge density of majority spin in Cr atom increases while that in S and Br atoms in the same layer decreases. The trend is opposite for spin minority. This indicates a spin transfer from S and Br to Cr, increasing the magnetic moment of the Cr atoms. In addition, we note that the reduction of spin in intralayer S and Br atoms is smaller than the increase of spin in Cr atom, suggesting that there is interlayer spin transfer (see Figure S9 in Supplemental Material [56]). This is in contrast to the spin transfer in metallic materials with both magnetic and nonmagnetic atoms where the spin transfer typically induces demagenetization where the magnetic momentum of magnetic atoms reduces [10]. Here we show that in the antiferromagnetic semiconductor, the optical laser pulsed induced inter-site spin transfer can also increase the magnetic moment of magnetic atoms. This mechanism is distinct from the recent study in ferromagnetic van der Waals monolayers where the SOC is necessary for the magnetization enhancement [46].
@@
When the light frequency is above the band gap, however, demagnetization of the Cr atoms happens as shown in Fig. 1(d). Figs. 1(e) and 1(f) show that the averaged atomic spin density of S and Br decreases, whereas the spin density of Cr first increases and then reduces to amounts less than the equilibrium in the majority spin channel. In addition, the intralayer spin increase of Br and S is much weaker than the spin decrease in Cr atom, indicating that the interlayer spin transfer dominates the demagnetization. We also computed the charge density in the interstitial region, as shown in the insets of Figs. 1(c) and 1(f) for below- and above-gap frequencies, respectively. The laser pulse is found to reduce the electron number in the interstitial region, indicating a corresponding increase in local electrons, consistent with the net electron gain on Cr atoms. However, whether the spins of these electrons align parallel or antiparallel to the magnetization direction depends on the laser frequency.
High fluence excitation.— In the high fluence cases (mJ/cm2), we find strong demagnetization that shows weak dependence on the laser frequency. We present the case with below-gap frequency and similar results for the above-gap case are presented in Supplemental Material [56]. Although the laser pulse does not induce a net magnetization due to the two-fold rotation symmetry about the -axis, it causes a strong reduction of the Néel vector. Fig. 2(a) shows the magnetic moment of a Cr atom in one magnetic layer, with the opposite moment in the other layer. As the laser pulse interacts with the system, the magnetic moment of Cr atoms is significantly reduced, directly contributing to the suppression of the Néel vector. Fig. 2(c) shows the spin-resolved charge transfer. The numbers of spin-majority and spin-minority electrons in Cr atoms change in a nearly symmetric fashion, indicating a spin-like transfer, i.e., net spin transfer dominates over net charge transfer. For nonmagnetic atoms (S and Br), the net magnetic moments show a slight initial enhancement and then decay to negligible values. These atoms show a charge-like transfer. As shown in Fig. 2(d), in S and Br atoms, both spin-majority and spin-minority electron numbers decrease, leading to a net loss of charge with relatively smaller changes in spin. This charge-like transfer in S and Br atoms generates free carriers in the interstitial region as shown in Fig. 2(b).
It is noted that the demagnetization shows a weak dependence on the SOC. In the absence of SOC, the spin-flip scattering is forbidden. As the intralayer charge transfer preserves the total magnetic moment, the strong reduction of magnetic moment in one magnetic layer, about 50%, must be attributed to the interlayer spin transfer. The presence of SOC does not show qualitative influence. The SOC starts to manifest itself near the end of the laser pulse, which introduces a weak correction of the demagnetization afterward and does not change the qualitative behavior. The demagnetization in one magnetic layer also shows weak dependence on the light frequency. We find similar qualitative behavior with above-gap frequency [56]. Quantitatively, higher demagnetization in Cr atoms whereas magnetization increase in S and Br are observed with higher frequency.
Ultrafast magnetization reorientation with magnetic fields.— The presence of a magnetic field profoundly alters the ultrafast magnetization dynamics. In the collinear spin configuration without a field, the laser pulse induces changes in the magnetization along the equilibrium direction. By contrast, an applied magnetic field enables an ultrafast reorientation of the magnetization vector, generating finite components perpendicular to the equilibrium direction. This transverse response modifies the relative spin angles on a sub-picosecond timescale, enabling an ultrafast initiation of spin dynamics and control of electronic structure and excitations coupled to the spin configuration.
We begin with the low-fluence regime. The magnetization dynamics along the easy axis with a magnetic field shows behavior similar to the zero-field case (see Fig. S7 in Supplemental Material [56]). Below we focus on the transverse response. At zero magnetic field , the net transverse components in one magnetic layer are zeros. In the unit cell of one layer, there are two Cr atoms (Cr1 and Cr2). Fig. 3(a) shows the transverse magnetic moment of Cr1 that show fast oscillation. Nevertheless, the total magnetic moment of Cr1 and Cr2 vanishes. The presence of a magnetic field alters the equilibrium spin configuration, inducing spin canting in - plane as shown in Fig. 3(b). The presence of a laser pulse changes the components of magnetic moment in and directions. As shown in Fig. 3(b), the net magnetization along direction in Cr1-Cr2 layer is opposite to that in Cr3-Cr4 layer. In contrast, the magnetic moment components along direction are the same for both layers. Such a spin configuration is invariant under a two-fold rotation operation around -axis, along which the magnetic field is applied. This configuration analogies the presence of coherent magnons that are even under the operation, i.e., the so called bright magnons in CrSBr. While the magnon frequency is in the order of 30 GHz [31, 32], the laser pulse induces a spin reorientation in 100 fs.
Then we turn to the high-fluence cases. The presence of a magnetic field suppresses the laser induced demagnetization while enhances the magnetization reorientation. Fig. 4(a) shows the net magnetic moment amplitude in magnetic layer 1 (L1). After the laser pulse, it reduces by about to that is much smaller than the demagnetization shown in Fig. 2(a). This is because the demagnetization is dominant by the interlayer hopping induced spin transfer as discussed in the previous section whereas the effective interlayer hopping is suppressed by the spin canting [57].
Figure 4(b) shows that a high-fluence laser pulse can effectively change the orientation of the magnetization vector in the presence of a magnetic field. This figure shows the time-evolution of the azimuthal and polar angles of the magnetization in L1. In the absence of the laser pulse, the magnetization lies in - plane with an azimuthal angle . The polar angle indicates that the spin deviates away from the easy axis due to the magnetic field. After the laser pulse, the azimuthal angle changes by about 10 degree at within 150 fs that continuously increases. This indicates a significant change in the spin configuration in an ultrafast time scale. The magnetization in layer 2 (L2) also changes. As shown in Fig. 4(c), the changes of magnetization in L1 and L2 along -direction are opposite while that along -direction are the same. This induces an ultrafast change of the relative angle between the magnetization in L1 and L2 as shown in Fig. 4(d). Such a spin configuration also corresponds to the generation of large amplitude coherent magnons in this system.
The ultrafast generation of coherent magnons challenges the often-invoked analogy with coherent phonons. Phonons, which are quantized lattice vibrations, originate from collective displacements of heavy nuclei and are thus inherently limited by the inertia of the lattice, with dynamics constrained by their characteristic vibrational frequencies [58, 59, 60]. In contrast, magnons are quantized spin waves, arising from collective reorientations of electronic spins. Since spins are emergent electronic degrees of freedom, they can be reconfigured almost instantaneously through ultrafast changes in electronic states, enabling responses on femtosecond timescales that is much faster than the intrinsic dynamics of magnetization. This fundamental distinction underscores that coherent magnon excitations are not simply magnetic analogues of phonons, but rather represent a unique pathway for ultrafast spin manipulation beyond the mechanically limited dynamics of the lattice.
Summary.—We studied the ultrafast dynamics of semiconducting antiferromagnetic CrSBr using the rt-TDDFT formalism. In the absence of a magnetic field, laser pulses only change the magnetization along the easy axis direction while the transverse components are not excited. We find that a low-fluence laser pulse with below-gap frequency can enhance the local magnetic moment and thus increase the Néel vector due to the spin transfer from nonmagnetic to magnetic layers. A high-fluence laser pulse, however, induces significant demagnetization in each magnetic layer and thus reduces the Néel vector by interlayer spin transfer between magnetic atoms. The nonmagnetic atoms, i.e., S and Br, however, show charge-like transfer that has weak contribution to demagnetization whereas generates free carriers.
The presence of a magnetic field significantly changes the ultrafast magnetization dynamics. Beside the change of the magnetitude of the magnetization, the magnetic field enables the ultrafast reorientation of the magnetization vector. The reoriented magnetizations in neighboring magnetic layers respect the two-fold rotation about -axis, corresponding to the generation of coherent magnons on the optical-magnon mode that is even under this symmetry operation. This mode changes the relative angle between neighboring magnetic layers, enabling periodic modulation of the electronic properties. These results deepen our understanding of how the coherent magnons are excited in van der Waals magnets and highlight the fundamental difference between coherent excitations of magnons and phonons, offering a solid starting point for future study of the subsequent magnetization dynamics.